# Study of quantum phase transition and entanglement in coupled top systems with standard and nonstandard symmetries under Floquet formalism

### v1

@@ II Model
The first two terms represent the Hamiltonian of the individual top [3], and the third term is the coupling between the two tops. This Hamiltonian is represented in terms of the angular momentum operators $\textbf{J}=(J_{xi},J_{yi},J_{zi})$ which follow usual angular momentum algebra $[J_{\alpha i},J_{\beta i}]=\imath\,\epsilon_{{\alpha\beta\gamma},i}\,J_{\gamma i}$ and $i=1,2$ denotes individual top. Here, $\imath=\sqrt{-1}$ and the standard Levi-Civita symbol $\epsilon$ is introduced to indicate the cyclic permutation in the commutator relations. As usual, the angular momentum operators corresponding to different tops commute, i.e., $[J_{\alpha 1},J_{\beta 2}]=0$, for $\alpha,\beta\in(x,y,z)$. We set the same spin size $j$ for both tops. The first term in the Hamiltonian describes the free precession of the individual top around $x$-axis with angular velocity $p_{i}/T$, and the second term describes torsion in the individual top about $z$-axis by an angle proportional to $J_{zi}$, where the proportionality constant $k_{i}$ is a dimensionless quantity. The second term is acting on the system in a periodic $\delta$-kicked fashion with time-period $T$, and it is responsible for the chaos in the individual top. The third term shows spin-spin interaction between the two tops, which also influences the system in a time-periodic $\delta$-kicked fashion with the same period $T$. The parameter $\varepsilon_{0}$ determines the interaction or coupling strength between them. The corresponding Floquet operator or the time-evolution operator between two consecutive kicks is:
$\begin{split}U(T)=&\,e^{-i\frac{\varepsilon_{0}}{j}J_{z1}\otimes J_{z2}}\left[e^{-i\frac{k_{1}}{2j}J_{z1}^{2}}\,e^{-ip_{1}J_{x1}}\right.\\ \otimes&\left.e^{-i\frac{k_{2}}{2j}J_{z2}^{2}}\,e^{-ip_{2}J_{x2}}\right].\end{split}$ (2)
III The effective Hamiltonian of the CKT system
We consider the driving frequency $\omega=2\pi/T$ much larger than other energy scales of the CKT model. Therefore, we employ the Van Vleck based perturbation method of Floquet formalism to derive the Floquet Hamiltonian, which is the effective time-independent Hamiltonian of the system. The general form of the effective Hamiltonian of any kicked system at the high-frequency limit up to $\mathcal{O}(\omega^{-2})$ is [4]:
$H_{\rm eff}=H_{0}+\frac{V}{T}+\frac{1}{24}\left[\left[V,H_{0}\right],V\right]$
$=H_{0}+\frac{V}{T}+\frac{1}{12}VH_{0}V-\frac{1}{24}\left(H_{0}V^{2}+V^{2}H_{0}\right),$ (3)
where $H_{0}$ is the static or undriven part of the Hamiltonian and $V$ is the Dirac $\delta$ kicked term. The third term of the above Hamiltonian appears from $\omega^{-2}$ order the perturbation, and this term is $\omega$-independent because we have substitute $\omega T=2\pi$. The presence of the factor of $1/12$ (or $1/24$) reduces the effect of the $\mathcal{O}(\omega^{-2})$ term on the classical dynamics as compared to the other terms. This also reflects in the quantum spectrum of the effective Hamiltonian $H_{\rm eff}$. Therefore, to simplify the analysis, we ignore the second-order terms. Now, the effective Hamiltonian becomes:
$H_{\rm eff}$ $=H_{0}+\frac{V}{T}=\frac{1}{T}\Bigl{[}\bigl{(}p_{1}J_{x1}+p_{2}J_{x2}\bigr{)}$
$+\frac{1}{2j}\bigl{(}k_{1}J_{z1}^{2}\otimes\mathbbm{1}+k_{2}\mathbbm{1}\otimes J_{z2}^{2}+2\varepsilon_{0}J_{z1}\otimes J_{z2}\bigr{)}\Bigr{]}.$ (4)
According to the definition given in the original Hamiltonian, the parameters $p_{i}$’s are the rotation angle, $k_{i}$s are the angle of torsion and $\varepsilon$ coupling within the time interval $T$. We now rescale these parameters as $\Omega_{i}=p_{i}/T,\,\kappa_{i}=k_{i}/T$, and $\varepsilon=\varepsilon_{0}/T$, where $\Omega_{i}$’s are the angular rotation rate or angular velocity, $\kappa_{i}$’s are the torsional rate, and $\varepsilon$ rate of coupling. Therefore, in terms of the rescaled parameters, the effective Hamiltonian becomes
$H_{\rm eff}$ $=\bigl{(}\Omega_{1}J_{x1}+\Omega_{2}J_{x2}\bigr{)}$
$+\frac{1}{2j}\bigl{(}\kappa_{1}J_{z1}^{2}+\kappa_{2}J_{z2}^{2}+2\,\varepsilon J_{z1}J_{z2}\bigr{)}.$ (5)
We have dropped the tensor product notations $\otimes$ in the above and will not use them in the remaining part of the paper. The suffixes $i=1,2$ denote the individual Hilbert space of the two tops, thus indicating the underlying tensor product structure of the two Hilbert spaces.
IV Classical dynamics of the effective Hamiltonian of the CKT
[…]

### v2

@@ II Model
The first two terms represent the Hamiltonian of the individual top [3], and the third is the coupling between the two tops. This Hamiltonian is represented in terms of the angular momentum operators $\textbf{J}=(J_{xi},J_{yi},J_{zi})$ which follow usual angular momentum algebra $[J_{\alpha i},J_{\beta i}]=i\epsilon_{{\alpha\beta\gamma},i}\,J_{\gamma i}$ and $i=1,2$ denotes individual top. The standard Levi-Civita symbol $\epsilon$ indicates the cyclic permutation in the commutator relations. As usual, the angular momentum operators corresponding to different tops commute, i.e., $[J_{\alpha 1},J_{\beta 2}]=0$, for $\alpha,\beta\in(x,y,z)$. We set the spin size of both the tops equal to $j$. The first term in the Hamiltonian describes the free precession of the individual top around $x$-axis with angular velocity $\Omega_{i}$, and the second term describes torsion in the individual top about $z$-axis by an angle proportional to $J_{zi}$, where the proportionality constant $\kappa_{i}$ is a dimensionless quantity. The third term describes the spin-spin coupling between the two tops of strength $\varepsilon$.
The classical Hamiltonian $H_{\rm{cl}}$ is obtained from Eq. (1) by dividing the quantum Hamiltonian by spin $j$ and set the limit $j\rightarrow\infty$. At the large spin limit, the rescaled angular momentum operators $J_{\alpha i}/j$ commute with each other, and we obtain the classical Hamiltonian as:
$H_{\rm{cl}}=\Omega_{1}X_{1}+\frac{1}{2}\kappa_{1}Z_{1}^{2}+\Omega_{2}X_{2}+\frac{1}{2}\kappa_{2}Z_{2}^{2}+\varepsilon Z_{1}Z_{2}.$ (2)
Here, $\alpha_{i}=J_{\alpha i}/j$, where $\alpha\in(X,Y,Z)$ denote different componets of the angular momentum and $i=1,2$ represents first and second tops.
III Classical dynamics of the coupled top Hamiltonian
From the classical Hamiltonian given in Eq. (2), we can derive classical equations of motion of the coupled top system using the generalized Poisson brackets introduced in Ref. [39]. The same equations of motion can also be obtained as follows: first, we write Heisenberg’s equation for each component of the angular momentum operator, i.e., $\frac{dJ_{\alpha i}}{dt}=-i\left[J_{\alpha i},H\right]$. Then divide both sides of the equation by the spin $j$. Finally, we use the commutation property of the rescaled angular momentum operators and obtain the equations of motion. The straightforwardness of the latter method prompts us to apply it to derive the classical equation of motion, and the following equations of motion are obtained (a detailed derivation is given in Appendix A)
$\dot{X}_{1}$ $=-Y_{1}\bigl{(}\kappa_{1}\,Z_{1}+\varepsilon\,Z_{2}\bigr{)},$
$\dot{Z}_{1}$ $=\Omega_{1}\,Y_{1},$
$\dot{X}_{2}$ $=-Y_{2}\bigl{(}\kappa_{2}\,Z_{2}+\varepsilon\,Z_{1}\bigr{)},$
$\dot{Z}_{2}$ $=\Omega_{2}\,Y_{2}.$ (3)
The rescaled angular momentum variables satisfy the constraint $X_{1}^{2}+Y_{1}^{2}+Z_{1}^{2}=X_{2}^{2}+Y_{2}^{2}+Z_{2}^{2}=1$. This suggests that the classical dynamics of the coupled top is restricted to the surface of $2$-spheres of unit radius. We can also exploit these constraints to reduce the degrees of freedom from six to four by transforming the Cartesian coordinates into spherical polar coordinates with unit radius. That is, we set the angular momentum variables as $X_{i}=\sin\theta_{i}\,\cos\phi_{i},\,Y_{i}=\sin\theta_{i}\,\sin\phi_{i},$ and $Z_{i}=\cos\theta_{i}$, where $Z_{i}$ and $\phi_{i}$ become canonically conjugate variables $\{Z_{i},\phi_{i}\}=1$ for the $i$-th top. In terms of these new variables $(\theta_{i},\phi_{i})$, the classical Hamiltonian becomes
$H_{\rm cl}$ $=\Omega_{1}\,\sqrt{1-Z_{1}^{2}}\,\cos\phi_{1}+\Omega_{2}\,\sqrt{1-Z_{2}^{2}}\,\cos\phi_{2}$
and the corresponding Hamilton’s equations of motion become:
For the torsion-free FP model, we set $\kappa_{1}=\kappa_{2}=0$ in the above set of equations. On the other hand, for the NZT-I model, we consider $\kappa_{1}=\kappa_{2}$; and for the NZT-II model, $\kappa_{1}=-\kappa_{2}$, where $\kappa_{i}$ with $i=1,2$ denote the strength of the torsional term of the $i$-th top. The symmetry properties of these tops are discussed extensively in Sec. IV.
III.1 FP model (torsion-free case)
Figure 1: The dynamics of the coupled top is projected on the phase space of the first top at the FP limit. The FP limit is obtained by setting $\kappa_{1}=\kappa_{2}=0$. We set $\Omega_{1}=\Omega_{2}=1.0$. Left panel: Coupling $\varepsilon=0.8$; and Right panel: $\varepsilon=1.3$.
Figure 1 presents the phase space dynamics of the coupled top at the FP limit. Here, we fix $\Omega_{1}=\Omega_{2}=1.0$ and consider two different coupling strengths $\varepsilon=0.8$ (Left panel) and $\varepsilon=1.3$ (Right panel). We have chosen these two coupling strengths because, at $\varepsilon=1.2$, the phase space dynamics of the FP system show a transition, which is clear from the appearance of substructures at the center.
III.2 Nonzero torsions case
We present the classical dynamics of the Hamiltonian $H_{\rm cl}$ of the coupled top. Here again, we set the parameters $\Omega_{1}=\Omega_{2}=1.0$ and consider the same two coupling cases, $\varepsilon=0.8$ and $1.3$, to show the effect of the nonlinear torsional terms. In Fig. 2, we consider two cases of torsions: at the top panels, two torsional terms are equal in magnitude and sign, i.e., $\kappa_{1}=\kappa_{2}$, and named this NZT-I model. At the bottom panels, we consider torsional terms with opposite signs, but of the same magnitude, i.e., $\kappa_{1}=-\kappa_{2}$, and named this NZT-II model. The phase space dynamics are qualitatively similar for the NZT-I model, presented in the top panels for both coupling strengths. This behavior is markedly different from the FP case, as discussed above. However, for the NZT-II model, the dynamics are very different from the case with the same value of the torsions. Here, we see that the trajectories are diffusing from one stable island to another at the center of the phase space of the first top. Here, we have not plotted the Poincaré section; instead, we have just projected the phase space trajectories of the coupled top residing on the four-dimensional phase space to the phase space of the first top having dimension two. Therefore, this diffusion of trajectories from one island to another happens through a higher-dimensional phase space. We also see the same behavior in other parts of the phase space.
III.3 Lyapunov Exponent for nonzero torsions case
In this section, we have studied the Lyapunov exponent $\Lambda$ as a measure of chaos of the coupled top model. This measure is defined as
$\Lambda=\lim_{t\rightarrow\infty}\lim_{\Delta(0)\rightarrow 0}\frac{1}{t}\ln\left[\frac{\Delta(t)}{\Delta(0)}\right].$ (6)
Here,
$\Delta(t)=\sqrt{\sum_{i=1}^{2}\Delta X_{i}(t)^{2}+\Delta Y_{i}(t)^{2}+\Delta Z_{i}(t)^{2}}$
is the separation between two trajectories at time $t$, which were initially separated by a tiny amount $\Delta(0)\rightarrow 0$. The Lyapunov exponent $\Lambda$ determines the exponential sensitivity of the initial condition of two neighboring trajectories. In Fig. 3, we have shown $\Lambda$ as a function of the coupling strength $\varepsilon$ for the non-zero torsion cases of the model. We have considered two initial conditions: (i) $Z_{1}=-0.7$ and $\Phi_{1}/\pi=0.1$ (shown as red circle in Fig. 2), and (ii) $Z_{1}=-0.1$ and $\Phi_{1}/\pi=0.1$ (shown as meganta circle in Fig. 2).
In Figs. 3(a) and (b), the results for NZT-I are presented for the above two initial conditions, whereas Figs. 3(c) and (d) present the results for NZT-II for the same initial conditions. For the NZT-I model, $\Lambda$ increases with the coupling strength $\varepsilon$ for both the initial conditions. However, for the initial condition marked by a red circle in Fig. 2, $\Lambda$ shows a peak $\varepsilon\simeq 0.3$; and then it goes down and saturates at a lower value when the coupling strength $\varepsilon\gtrsim 0.5$. For the other initial condition, we only observe a very slow increment in $\Lambda$ as a function of $\varepsilon$. In the NZT-II model, the Lyapunov exponent $\Lambda$ also shows variation with the coupling strengths $\varepsilon$. Moreover, the $\Lambda$ is greater than the NZT-I model for sufficiently strong coupling. This suggests that the NZT-II model is more chaotic. A transition in $\Lambda$ has been observed around $\varepsilon=1.0$. We will see later that this transition in classical dynamics influences the entanglement transition and the transition in the excited state energy of the systems. This behavior of $\Lambda$ can be understood better from the behavior of a bunch of phase space trajectories with initial points within the magenta and red circles mentioned above. For the NZT-I and NZT-II models, the phase space trajectories are shown for different values of the coupling strengths $\varepsilon$ in Figs. 7 and 8 of Appendix B.
Figure 3: Lyapunov exponent is plotted with the variation in coupling strength. Top row: For NZT-I ($\kappa_{1}=\kappa_{2}=1$) and Bottom row: For NZT-II ($\kappa_{1}=-\kappa_{2}=1$). Circles in the red and magenta colors in the figures show the Lyapunov exponent $\Lambda$, and the black solid line shows the fitted curve of the data.
IV Symmetry properties of the coupled top Hamiltonian
IV.1 FP model (torsion-free case)
For the FP model, the classical Hamiltonian of the coupled top becomes
$H_{\rm FP}=\Omega_{1}J_{x1}+\Omega_{2}J_{x2}+\frac{\varepsilon}{j}J_{z1}J_{z2}$ (7)
$U_{0}=e^{-i\pi J_{x1}}\otimes e^{-i\pi J_{x2}},$ (8)
$C=e^{i\alpha}e^{-i\pi J_{z1}}\otimes e^{-i\pi J_{y2}},$ (9)
which gives $C\,H_{\rm FP}\,C^{\dagger}=-H_{\rm FP}$ or $\left(C\,H_{\rm FP}+H_{\rm FP}\,C\right)=0$. This implies that ${\rm Tr}\,(H_{\rm FP})=0$. The phase factor $e^{i\alpha}$ is not playing any role here to show this symmetry of $H_{\rm FP}$ under $C$ because its complex conjugate factor from $C^{\dagger}$ part will trivially cancel it. However, this innocent-looking phase factor will become important while classifying the system under different symmetry classes. The operator, $C$, is called the chirality operator. In the case of the integer spin $C^{2}=\mathbbm{1}$. This property of $H_{\rm FP}$ implies the presence of nonstandard symmetries different from the standard Wigner-Dyson three-fold symmetry classes. In addition, the FP Hamiltonian has a nonunitary symmetry, like the time-reversal symmetry. We denote the time-reversal operator $\mathcal{T}$, which, on a standard basis, flips the sign of the $J_{y}$ operator but keeps the other two angular momenta invariant. Therefore, $H_{\rm FP}$ also satisfies $\mathcal{T}\,H_{\rm FP}\mathcal{T}^{-1}=H_{\rm FP}$. Depending on whether the chirality operator $C$ commutes with $\mathcal{T}$, we get different classes of symmetries for $H_{\rm FP}$. If the spin $j$ is integer, then we set $\alpha=0$ and find that
$\mathcal{T}\,C\,\mathcal{T}^{-1}$ $=\mathcal{T}\left[e^{-i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}\right]\mathcal{T}^{-1}$
$=e^{i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}$
$=e^{2i\pi J_{z1}}\,e^{-i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}$
$=C$ (10)
where we use $\mathcal{T}\,i\,\mathcal{T}^{-1}=-i$; and for integer $j$, the term $e^{2i\pi J_{z1}}$ becomes identity matrix. Thus we find $\mathcal{T}\,C\,\mathcal{T}^{-1}=C$. However, for the half-integer spin $e^{2i\pi J_{z1}}\neq\mathbbm{1}$ and hence $\mathcal{T}\,C\,\mathcal{T}^{-1}\neq C$. For this case, the phase $\alpha$ can be tuned to get a $C$ which commutes with $\mathcal{T}$. For any arbitrary phase $\alpha$, we have,
$\mathcal{T}\,C\,\mathcal{T}^{-1}$ $=\mathcal{T}\left[e^{i\alpha}e^{-i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}\right]\mathcal{T}^{-1}$
$=e^{-i\alpha}e^{i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}$
$=e^{-i2\alpha}\,e^{i2\pi J_{z1}}\,\left[e^{i\alpha}\,e^{-i\pi J_{z1}}\,\otimes e^{-i\pi J_{y2}}\right]$
$=e^{i2(\pi J_{z1}-\alpha)}\,C.$ (11)
$C=e^{i\pi j}e^{-i\pi J_{z1}}\otimes e^{-i\pi J_{y2}}$ (12)
IV.2 Nonzero torsions case
For the nonzero torsions, first consider the NZT-I model ($\kappa_{1}=\kappa_{2}=\kappa$) and $\Omega_{1}=\Omega_{2}=\Omega$. Then the classical Hamiltonian $H_{\rm CT}$ of the coupled top becomes:
$=H_{\rm FP}+\frac{\kappa}{2j}\left(J_{z1}^{2}+J_{z2}^{2}\right)$
$\equiv H_{\rm FP}+H_{\rm NL},$ (13)
$=\frac{1}{j}\sum_{m=1}^{j}m^{2}\left(\kappa_{1}+\kappa_{2}\right)$
$=\frac{1}{6}(j+1)(2j+1)\left(\kappa_{1}+\kappa_{2}\right).$ (14)
The above relation shows that the trace is zero when $\kappa_{1}=-\kappa_{2}=\kappa$. Therefore, the coupled top Hamiltonian with trace zero is of the form
We have observed very different classical dynamics for this particular case of torsion with opposite signs. However, the Hamiltonian $H_{\rm CT}$ still lacks chiral symmetry under the transformation of $C$. This result implies that there exists a different chirality operator, say $C^{\prime}$, under which $C^{\prime}\,H_{\rm CT}\,C^{\prime\dagger}=-H_{\rm CT}$. One point is to be noted that now the Hamiltonian $H_{\rm CT}$ is not symmetric under permutation. We exploit this fact and obtain $C^{\prime}=PC$, where $P$ is the permutation operator. One can now easily check that $C^{\prime}\,H_{\rm CT}\,C^{\prime\dagger}=P\,C\,H_{\rm CT}\,C^{\dagger}P^{\dagger}=-H_{\rm CT}$ or $\left(C^{\prime}\,H_{\rm CT}+H_{\rm CT}\,C^{\prime}\right)=0$. Since the chirality operator $C$ is time-reversal symmetric, one can show that the chirality operator $C^{\prime}$ for the coupled top Hamiltonian is also time-reversal symmetric. Moreover, for $C^{2}=\pm\mathbbm{1}$, the other chirality operator also satisfies the same property, i.e., $C^{\prime 2}=\pm\mathbbm{1}$. Therefore, the coupled top Hamiltonian can also be classified into two chiral symmetry classes, BDI and CI [30, 24]. Besides, for the nonzero trace case, one can not find a chirality operator under which the coupled top Hamiltonian will show the symmetry property. Therefore, this Hamiltonian belongs to the standard symmetry class.
V Entanglement Calculation
We now study the entanglement property of the coupled-top system represented by the Hamiltonian $H_{\rm CT}$ for zero (FP case) and nonzero torsion cases (NZT-I and NZT-II models). We study the entanglement of the ground and the most excited states of the coupled top Hamiltonian. Since these are pure states, we use the von Neumann entropy $S_{V}$ of the reduced density matrix (RDM) corresponding to these states as a measure of entanglement. The RDM corresponding to one of the tops is obtained by taking the partial trace over the other. If $|\psi\rangle$ is a state of the coupled top, then the RDM corresponding to one of the tops will be $\rho_{i}={\rm Tr}_{\overline{i}}\left(|\psi\rangle\langle\psi|\right)$, where $(i,\overline{i})=1,2\,(\overline{i}=2,1)$ represents two tops. Then the von Neumann entropy is defined as follows:
$S_{v}=-{\rm Tr}_{1}[\rho_{1}\,\ln\rho_{1}]=-{\rm Tr}_{2}[\rho_{2}\,\ln\rho_{2}]=-\sum_{i}\lambda_{i}\ln\lambda_{i}.$ (16)
Here, $\rho_{1}$ and $\rho_{2}$ share the same nonzero eigenspectrum $\{\lambda_{i}\}$.
     Ground State            Excited State
Figure 4: Showing the variation of entanglement (von Neumann entropy) with coupling strength. The left column presents the results for the ground state, and the right column presents the results for the most excited state.
Figure 4 presents the results of the entanglement calculation. Here, the variation of the von Neumann entropy is shown as a function of the coupling strength between the two tops $\varepsilon$ for the spin $j=10$. We show the entanglement of two states: the ground state and the most excited state. Figure 4(a) shows the result for the ground state energy of the FP model (i.e., $\kappa_{1}=\kappa_{2}=0$). We find that, initially, $S_{V}$ increases linearly with $\varepsilon$ and attains a maximum value $S_{V}^{\rm max}\simeq 0.91$ at $\varepsilon=1.2$, and it then immediately starts falling with the increment of the coupling strength. Later, we show that this transition in entanglement is consistent with the phase transition in the system. The entanglement falls till the coupling strength $\varepsilon\simeq 3.0$ and saturates at a value close to $\ln 2$ with some fluctuations. This saturation of $S_{V}$ around $\ln 2$ indicates permutation symmetry in the system, i.e., if we interchange the two tops, the Hamiltonian remains invariant. If we consider two non-identical coupled tops by setting unequal $\Omega$, the permutation symmetry will be broken, and consequently, the entanglement between the tops approaches zero for higher coupling strengths. In Fig. 4(b), we present the results for the most excited state. Here also, the transition in the entanglement happens at the same coupling strength $\varepsilon=1.2$. Interestingly, we do not see the effect of permutation symmetry on the excited state, and it becomes completely unentangled with $S_{V}=0$ for stronger coupling strengths.
In Figs. 4(c) and (d), the von Neumann entropy $S_{V}$ is presented for the NZT-I model. In this case, like the previous one, the entanglement of the ground state initially increases linearly with the coupling strength $\varepsilon$ and reaches a maximum $S_{V}^{\rm max}\simeq 1.0$ at $\varepsilon=2.0$. The entanglement then decreases and finally saturates at $\ln 2$ with some fluctuations for stronger coupling strengths $\varepsilon\gtrsim 4.0$. This saturation of entanglement at $S_{V}\simeq\ln 2$, also observed in the FP model, can be understood in the following way. At the strong coupling limit, the individual dynamics of the tops become almost negligible. The spin-spin coupling term in the Hamiltonian becomes the dominant term, and then the energy eigenstates approximately become $|m_{1}\rangle\otimes|m_{2}\rangle$, where $|m_{i}\rangle$’s are the eigenstates of $J_{z_{i}}$ and $m_{i}=-j,\dots,+j$. These eigenstates are unentangled states. However, due to the presence of permutation symmetry in the system, $|\psi_{m_{1},m_{2}}\rangle=\frac{1}{\sqrt{2}}\left(|m_{1}\rangle\otimes|m_{2}\rangle\pm|-m_{1}\rangle\otimes|-m_{2}\rangle\right)$ are also valid eigenstates with the same energy. The von Neumann entropy of these eigenstates $|\psi_{m_{1},m_{2}}\rangle$ equals to $\ln 2$. However, this nonzero entanglement of these states is useless because it cannot be used as a resource for any quantum protocols. This is just an artifact of the underlying permutation symmetry in the system.
Figures 4(e) and (f) show the results for the NZT-II model. This version of the coupled top system is interesting because it follows one of the nonstandard symmetry classes [30, 24]. Since we consider integer spin here, the system follows the BDI symmetry class. In this case, the ground state entanglement also increases linearly and reaches its maximum $S_{V}^{\rm max}\simeq 0.8$ at $\varepsilon\simeq 1.0$ and suddenly starts decreasing. Here, the entanglement transition point is the same as the FP model. Like the previous two cases, the entanglement immediately goes down to $\sim\ln 2$ immediately after reaching the maximum, and then for $\varepsilon\simeq 3.0$, the system becomes unentangled. Interestingly, for $\varepsilon\gtrsim 3.0$, the von Neumann entropy increases slowly in a linear fashion with some fluctuations. The behavior of the excited state entanglement is almost similar to the ground state, except we observe that the excited state becomes unentangled or a product state at $\varepsilon\simeq 3.0$ and remains as such for larger values of coupling strength.
Overall, we find that the entanglement transition in the ground state of the FP model and the NZT-II model takes place at the same value of the coupling strength $\varepsilon\simeq 1.0$. However, the entanglement of the ground state at the transition point is higher for the FP model. Notably, the FP model has permutation symmetry, but the NZT-II model does not have that symmetry. Therefore, the saturation values of the entanglement for these two models are different: the FP model shows spurious nonzero entanglement with $S_{V}\sim\ln 2$, and the other model becomes completely unentangled and then increases slowly with the coupling strength. The entanglement of the most excited state behaves qualitatively similarly for both these models. Now, for the NZT-I model, the ground state entanglement makes a transition at a larger coupling strength $\varepsilon\simeq 2.0$ and then the entanglement reaches the saturation value $S_{V}=\ln 2$ around $\varepsilon\simeq 4.0$. The entanglement of the excited state of this model transitions at a very small value of coupling strength $\varepsilon=0.3$, after that the entanglement decreases till $\varepsilon=2.0$. The model transitions to the unentangled state in the $\varepsilon>2.0$ regime.
VI Transition in energies and correlation with entanglement
We now investigate the phase transition in the system by studying the variation in the FP model’s excited state and ground state energies and the coupled top models with non-zero torsion cases. We find that these transitions are related to the transition in the entanglement between the two tops. We analyze all the results from the underlying classical dynamics. Notably, we have shown how the system’s stable-to-unstable transition of the classical fixed point (CFP) influences the quantum transitions. We observe that the quantum transitions and the transition in the classical fixed points (CFPs) stability occur at the same coupling strength $\varepsilon$ for torsion-free and nonzero torsional models.
The CFPs are determined by setting the time-derivatives of the dynamical variables $(Z_{i},\phi_{i})$ for $i=1,2$ at the left-hand side of Eq. (5) equal to zero. We now obtain a set of four coupled homogeneous algebraic equations, and the solutions of these equations give the CFPs of the system for different coupling strengths $\varepsilon$. We find the same four CFPs for the torsion-free and nonzero torsional cases:
The generic behavior of the CFPs of the system is the following: Initially, for very weak coupling strengths $\varepsilon\gtrsim 0.0$, the system has two fixed points (CFP-I and CFP-II), when both the spins of the coupled top are in the same direction ($X$-direction) with zero magnetization along the $Z$-axis, as shown in Fig. 5. This figure schematically shows that CFP-I and CFP-II are distinguished from the direction of the spin of the individual top: for CFP-I, both spins are along the negative $X$-direction, whereas for the CFP-II case, the spins are along the positive $X$-direction. These two fixed points correspond to symmetry-unbroken stable fixed points. At some critical values of the coupling strength $\varepsilon=\varepsilon_{\rm c}$, the classical fixed points become unstable and transition from the CFP-I to the CFP-III by bifurcation in energies. However, the critical value $\varepsilon_{\rm c}$ differs for CFP-I and CFP-II. The CFP-I state bifurcates and forms a pair of symmetry-broken anti-ferromagnetic states, CFP-III, at a critical point. On the other hand, the CFP-II state bifurcates at a different critical point and forms a ferromagnetic state, CFP-IV. These transitions are shown by a schematic diagram in Fig. 5.
We now present a detailed analysis of the transitions in energies for the torsion-free and nonzero torsional cases of the coupled top. In the left column of Fig. 6, we show the variation of the ground and the excited state energies with the coupling strength $\varepsilon$ to analyze the transition in energies of the system. In the right column, we have shown the corresponding variation of the quantum entanglement between the two tops. However, here we focus on the transition in energies and its relation with the entanglement, so we have shown the entanglement for the coupling strengths $\varepsilon\in[0.0,3.0]$.
VI.1 FP model (torsion-free case)
In the FP model, Fig. 6(a) shows the ground and excited states energy variation with the coupling strength $\varepsilon$. Here, the bifurcation is observed in both states. The ground state is stable at $E=-2.0$ when $\varepsilon<1.2$ (CFP-I state). Then, this state transitions to the CFP-III state at $\varepsilon=1.2$, which remains stable for $\varepsilon>1.2$. In the CFP-III state, the fixed point varies with the coupling strength as the ground state energy varies with the coupling strength as $Z_{1}=-Z_{2}=\sqrt{1-\frac{1}{\varepsilon^{2}}}$ with $\phi_{1}=\phi_{2}=\pi$, and consequently, the ground state energy varies as $E=-(\varepsilon+\frac{1}{\varepsilon})$.
On the other hand, for $\varepsilon<1.2$, the excited state energy of the system is stable at $E=+2$ (CFP-II state). This state transitions to the CFP-IV state by bifurcation at the critical point $\varepsilon=1.2$, and then it remains stable for $\varepsilon>1.2$. After the transition of the excited state from CFP-II to CFP-IV state, the fixed point also varies with $\varepsilon$ in the same fashion as the CFP-III state of the ground state. However, the excited state energy now increases as $E=(\varepsilon+\frac{1}{\varepsilon})$.
For the FP model, we observe that the transitions in ground and excited states are happening around the same value of the coupling strength $\varepsilon\simeq 1.0$. Figure 6(b) shows the variation in entanglement (von Neumann entropy) of the ground state (in red color) and excited state (in blue color) as a function of the coupling strength. We find that the entanglement of the ground state and the excited state maintains a steady growth till and then suddenly falls at $\varepsilon\simeq 1.0$. Interestingly, this transition in the entanglement is concurrent with a transition in energies and the stability of the fixed points of the underlying classical dynamical system.
VI.2 Nonzero torsion case
We now present the results for two nonzero torsional coupled top models: (a) NZT-I model ($\kappa_{1}=\kappa_{2}=1.0$) and (b) NZT-II model ($\kappa_{1}=-\kappa_{2}=1.0$).
VI.2.1 NZT-I model ($\kappa_{1}=\kappa_{2}=1.0$)
Figure 6(c) shows the variation in ground state and excited state energy as a function of the coupling strength $\varepsilon$. This figure shows that the CFP-I state becomes unstable at $\varepsilon=(\kappa_{1}+1)=2$ and bifurcates into the CFP-III state. The energy of the CFP-III state varies with the coupling strength as $E=(1-\varepsilon)+\frac{1}{1-\varepsilon}$. On the other hand, the CFP-II state becomes unstable when $\varepsilon>(-\kappa_{1}+1)=0$. However, the effect of the unstable CFP-II state becomes prominent at a finite but very small value $\varepsilon\simeq 0.23$, where the CFP-II state bifurcates into the CFP-IV state. The energy of the CFP-IV state follows $E=(1+\varepsilon)+\frac{1}{(1+\varepsilon)}$. In Fig. 6(d), the transition in entanglement is again exactly happening at the point of bifurcation for the ground and excited states. However, unlike the FP model, the transitions in the ground and excited states do not coincide at the same coupling strength. Still, these transitions are consistent with the corresponding entanglement transition and the bifurcation of the steady states.
VI.2.2 NZT-II model ($\kappa_{1}=-\kappa_{2}=1.0$)
The results corresponding to this case are shown in Figs. 6(e) and (f). From Fig. 6(e), we observe the transitions in the ground and excited states simultaneously at $\varepsilon\simeq 1.0$. This result is similar to the FP model case. Here, at the classical limit, the effect of torsion is canceled because of the opposite sign with the same magnitude of the torsion strength of each top. The energy for the CFP-III is calculated as $E=-\frac{2}{1-\varepsilon}-\varepsilon\left[1-\frac{1}{(1-\varepsilon)^{2}}\right]$; whereas for the CFP-IV state, the energy follows the relation $E=\frac{2}{1+\varepsilon}-\varepsilon\left[1-\frac{1}{(1+\varepsilon)^{2}}\right]$. Figure 6(f) shows the transition in entanglement, which is consistent with the energy transition.
The above analysis reveals that the energy transition coincides around $\varepsilon\simeq 1.0$ in the FP and NZT-II models. A common feature of these two models is that they both have chiral symmetry and thus follow one of the nonstandard symmetry classes. Since we only consider integer spin here, these models follow the BDI symmetry class. On the other hand, the NZT-I model does not follow chiral symmetry. Interestingly, the transition in energies does not coincide in this model. However, these transitions are consistent with the corresponding entanglement transition and the transition in the stability of the fixed point of the underlying classical system.
VII Conclusion
[…]